OpenStudy (anonymous):

can someone help. i dont understand how to find the equation with the slope being 3/5 ??Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (campbell_st):

well the slope intercept form of the line is y =mx + b m = slope of the line and b = y- intercept in your question m = 3/5 so you have y = 3/5 x + b substitute the values for the point(2, -4) that is x = 2 and y = -4 to find b so -4 = 3/5(2) + b or -4 = 6/5 + b so the value of b is -26/5 the equation of the line is y = 3/5 x - 26/5 to simplify things multiply every term by 5 then 5y = 3x - 26 or rewriting in the general form 3x - 5y = 26 hope this makes sense